OpenStudy (anonymous):

solve. find three consecutive integers whose sum is the square of 6

OpenStudy (anonymous):

so the equation you want is: \[x + x + 1 + x +2 = 36\]

OpenStudy (anonymous):

you combine like terms to get: \[3x + 3 = 36\]

OpenStudy (anonymous):

subtract 3 to get:

OpenStudy (anonymous):

\[3x = 33\]

OpenStudy (anonymous):

divide by 3 to get: \[x = 11\]

OpenStudy (anonymous):

so the three numbers are 11, 11+1 (12), and 11 + 2 (13)

OpenStudy (anonymous):

and finally: \[11 + 12 + 13 = 36 = 6^2\]

OpenStudy (anonymous):

so it works

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

could you help with this one..the greater of two consecutive integers is 10 less than twice the smaller. find the integers.

OpenStudy (anonymous):

so: \[x + 1 = 2x - 10\]

OpenStudy (anonymous):

add 10 to both sides: \[x + 11 = 2x\]

OpenStudy (anonymous):

subtract x: \[x = 11\]

OpenStudy (anonymous):

check: \[2(11) = 22 = 12 + 10\]

OpenStudy (anonymous):

so the integers are 11 and 12

OpenStudy (anonymous):

i meant to put two consecutive even** integers

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

o.k.

OpenStudy (anonymous):

\[x + 2 = 2x - 10\]

OpenStudy (anonymous):

because you skip a number

OpenStudy (anonymous):

so you get x is 12 through the same process as above

OpenStudy (anonymous):

check: \[2(12) = 24 = 14 + 10\]

OpenStudy (anonymous):

it works, so the numbers are 12 and 14

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

you are welcome